# Density estimation with LLMs: a geometric investigation of in-context learning trajectories

### v2

@@ 5 Conclusions
Inspired by emergent, in-context abilities of LLMs to continue stochastic time series, the current work explores the efficacy of foundation models operating as kernel density estimators; a crucial element within such time series forecasting. Through an application of InPCA, we determine that such in-context kernel density estimation proceeds within a common, low dimensional probability space; along meaningful trajectories that allow for a comparison between histogram, Gaussian kernel density estimation, and LLM in-context kernel density estimation. Through the lens of InPCA, it becomes clear that a profitable characterization of this in-context learning can be made in terms of a two-parameter, adaptive kernel density estimation framework; hinting at mechanistic basis (Appendix A.1) that points in the direction of future research.
Future direction: towards dispersive induction heads. Recent research has identified induction heads as fundamental circuits underlying many in-context learning behaviors in stochastic systems with discrete states (Olsson et al., 2022; Bigelow et al., 2024). These emergent circuits increase the predicted probability of token combinations that are observed in-context. However, such discrete mechanisms are insufficient to explain the in-context learning of continuous stochastic systems, such as the density estimation task we’ve studied. To address this gap, our Bespoke KDE analysis in Section 4.3.2 reveals that LLaMA might possess a kernel-like induction mechanism, which we term dispersive induction head. This is an extension to the induction head concept, and operates as follows:
  • Unlike standard induction heads (Akyürek et al., 2024), a dispersive induction head increases the predicted probability of not just exact matches, but also of similar tokens or words.
@@ A.3 Bespoke KDE cannot imitate all DE processes
A.3.1 InPCA visualization of original DE trajectories vs. Bespoke KDE imitations
@@
Figure 10: Bespoke KDE parameters fitted to Bayesian histogram. The constant fitted kernel width likely reflects the fixed bin width, while the near-zero shape parameter indicates a highly peaked distribution. High uncertainties suggest a fundamental mismatch between the models.
@@
These results highlight that while Bespoke KDE can effectively model certain DE processes (e.g., LLaMA models and Gaussian KDE), it cannot capture fundamentally different approaches like the Bayesian histogram.
A.3.2 Meta-InPCA Embedding of Trajectories
@@
A.4 Comparison of InPCA and standard PCA
@@
A.5 Additional notes on Kernel Density Estimation
@@ A.5.1 Optimal Bandwidth
A.5.2 Optimal kernel shape
@@
A.5.3 MISE vs. valid statistical distances
@@
A.6 Adaptive Kernel Visualization: Comparing LLaMA-2 with KDE
@@
A.7 Randomly generated PDFs
@@
A.8 KDE with Silverman bandwidth schedule
@@
with $C=1$.However, as noted in Appendix A.5.1, in practice one often estimates the pre-coefficient from data, such as Silverman’s rule of thumb (Equation 14). In this section, we replicate Figures 3 and 4 from Section 4, and use Silverman’s rule-of-thumb bandwidth instead of $C=1$. As shown in Figures 35 and 36, even with Silverman’s pre-coefficient, Gaussian KDE still shows a clear bias towards the Gaussian submanifold.

### v3

@@ 5 Conclusions
Inspired by emergent, in-context abilities of LLMs to continue stochastic time series, the current work explores the efficacy of foundation models operating as kernel density estimators; a crucial element within such time series forecasting. Through an application of InPCA, we determine that such in-context kernel density estimation proceeds within a common, low-dimensional probability space; along meaningful trajectories that allow for a comparison between histogram, Gaussian kernel density estimation, and LLM in-context kernel density estimation. Through the lens of InPCA, it becomes clear that a profitable characterization of this in-context learning can be made in terms of a two-parameter, adaptive kernel density estimation framework; hinting at mechanistic basis (Appendix A.1) that points in the direction of future research.
Future direction: towards dispersive induction heads. Recent research has identified induction heads as fundamental circuits underlying many in-context learning behaviors in discrete-state stochastic systems (Olsson et al., 2022; Bigelow et al., 2024). These emergent circuits increase the predicted probability of token combinations that are observed in-context. However, such discrete mechanisms are insufficient to explain the in-context learning of continuous stochastic systems, such as the density estimation tasks we have studied. To address this gap, our Bespoke KDE analysis in Section 4.3.2 reveals that LLMs might possess a kernel-like induction mechanism, which we term dispersive induction head. This is an extension to the induction head concept, and operates as follows:
  • Unlike standard induction heads (Akyürek et al., 2024), a dispersive induction head increases the predicted probability of not only exact matches, but also similar tokens or words.
@@ A.3 Bespoke KDE cannot imitate all DE processes
A.3.1 DE trajectories of LLaMA-2 7B, 13B, and 70Bf vs. Bespoke KDE imitations
@@
Figure 10: Bespoke KDE parameters fitted to Bayesian histogram. The constant fitted kernel width likely reflects the fixed bin width, while the near-zero shape parameter indicates a highly peaked distribution. The high uncertainties suggest fundamental mismatch between the two models.
@@
These results highlight that while Bespoke KDE can effectively model certain DE processes (e.g., LLMs and Gaussian KDE), it cannot capture fundamentally different approaches like the Bayesian histogram.
A.3.2 Meta-InPCA Embedding of Trajectories: the LLaMA-2 suite
@@
A.4 Meta-InPCA Embedding of Trajectories: LLaMA, Gemma, and Mistral
This section documents additional experiments on three other recently released LLMs: Gemma-2b and -7b (GemmaTeam et al., 2024) and Mistral-7b-v0.3 (Jiang et al., 2023). These models share similar tokenizers as LLaMA-2, and work well with the Hierarchy-PDF algorithm Liu et al. (2024).
Figure 12: Comparison of in-context density estimation trajectories for LLaMA, Gemma, and Mistral. Top row: 2D InPCA embeddings of DE trajectories for narrow and wide Gaussian targets. Bottom row: Corresponding ground truth distributions. These 2D embeddings capture 91% of pairwise Hellinger distances between probability distributions.
As shown in Figure 12, the in-context DE trajectories of LLaMA-2, Mistral v0.3, and Gemma, are strikingly similar, despite the fact that they were built and trained entirely independently by different teams.
Figure 13: Meta-InPCA embedding of DE trajectories ($\triangle$) and their bespoke KDE imitations ($\circ$). This 2D embedding captures 91% of pairwise meta-distances between trajectories.
To quantify the similarity of these models’ learning trajectories, we use the meta-InPCA embedding introduced in Section A.3.2 to visualize the pairwise trajectory distance from these models. As shown in Figure 13, Gemma and Mistral are well-approximated by bespoke KDE counterparts. In fact, all kernel-like methods are quite similar to each other, forming a cluster on the right corner of the embedding, with Bayesian histogram isolated on the left, similar to Figure 11.
Interestingly, Gemma-2b seems more similar to Mistral-7b-v0.3 and LlaMA-2 70b, than to Gemma-7b. Models from the same suites are not always the most alike in terms of in-context DE trajectories.
To summarize, the kernel-like density estimation process observed in this work is not limited to any specific suites of LLMs. We therefore speculate that LLMs might spontaneously converge to a universal in-context DE algorithm.
A.5 Appendix: Comparison of InPCA and standard PCA
@@
A.6 InPCA embeddings with symmetrized KL-divergence and Bhattacharyya distance
Figure 15: InPCA embeddings of LLaMA-2 in-context DE trajectories using symmetrized KL-divergence and Bhattacharyya distance. The target function is narrow Gaussian distribution as shown in Figure 2. Red dots denote negative eigenvalue, and red axis denotes imaginary principal component. (a) 2D InPCA embedding preserves 90% of pairwise Bhattacharyya distances. (b) 3D InPCA embedding preserves 87% of pairwise symmetrized KL-divergence.
Bhattacharyya distances (Quinn et al., 2019) symmetrized KL-divergence (Teoh et al., 2020) are two other statistical divergences commonly used for InPCA embeddings. They are defined as:
$D^{2}_{\text{BT}}(p,q)$ $=-\ln\left(\int\sqrt{p(x)q(x)}\,\textup{d}x\right)$ (11)
$D^{2}_{\text{sKL}}(p,q)$ $=\int p(x)\ln\frac{p(x)}{q(x)}\,\textup{d}x+\int q(x)\ln\frac{q(x)}{p(x)}\,\textup{d}x$ (12)
Figure 15 shows the InPCA embeddings resulting from these two distances. Both of these distances locally agrees with the Hellinger distance as well as the KL-divergence (Teoh et al., 2020). In fact, the Bhattacharyya distance has a deep physical connection to Hellinger distance, which was discussed in Quinn et al. (2019).
However, unlike the Hellinger distance used in this work, both Bhattacharyya distance and symmetrized KL-divergence violate the triangle in-equality (Quinn et al., 2019), and are therefore not valid distance metrics. Consequently, the eigenvalue decomposition step (Section 3.1) might result in negative eigenvalues, with the corresponding coordinates being imaginary. A linear space with a mixture of real and imaginary axes is a generalization of the Minkowski space, which is useful in special relativity (Einstein, 1905), where events are represented as points in spacetime with one timelike (imaginary) dimension, and up to three spacelike (real) dimensions. The Minkowski metric between two points $p=(x,t)$ and $q=(x^{\prime},t^{\prime})$ is given by:
$d^{2}(p,q)=(x-x^{\prime})^{2}-(t-t^{\prime})^{2}$ (13)
In our visualization context, the imaginary axis (colored red in Figure 15) plays a role similar to the time dimension in Minkowski space. The further separated two points are along the imaginary axis, the closer they are in real distance.
As shown in Panel (a), visualization using the Bhattacharyya distance is quantitatively the same as its Hellinger counterpart, shown in Figure 2. Visualization with symmetrized KL-divergence, shown in the top figure in Panel (b), is also qualitatively similar, except that here LLaMA’s estimated density no longer seems to converge to the ground truth (blue dot). However, this is a false impression. The bottom figure in Panel (b) reveals a third principal component which is imaginary. The end point of LLaMA’s estimated density is separated from the ground truth (blue dot) along both the real and imaginary axes, by about the same distance. The positive and negative distances from spacelike and timelike seprations roughly cancel out, which means the statistical distance between LLaMA’s estimated density and the ground truth is actually small.
Negative eigenvalues from these metrics might be physically meaningful in certain circumstance, such as shown in Teoh et al. (2020); Quinn et al. (2019); Mao et al. (2024). However, for the purpose of visualizing in-context DE trajectories, they complicate the geometric analysis without adding much insight.
A.7 Additional notes on Kernel Density Estimation
@@ A.7.1 Optimal Bandwidth
A.7.2 Optimal kernel shape
@@
A.7.3 MISE vs. valid statistical distances
@@
A.8 Adaptive Kernel Visualization: Comparing LLaMA-2 with KDE
@@
A.9 Randomly generated PDFs
@@
A.10 Heavy-tailed distributions
In this section, we visualize the in-context DE trajectories of Student’s t-distribution (Ahsanullah et al., 2014). The t-distribution is parametrized as
$p_{v}(x)=\frac{\Gamma(\frac{v+1}{2})}{\sqrt{v\pi}\Gamma(\frac{v}{2})}\left(1+\frac{x^{2}}{v}\right)^{-\frac{v+1}{2}}$ (21)
where $v$ controls the amount of probability mass in the tails. Smaller $v$ leads to heavier tails. For $v=1$, the t-distribution reduces to the Cauchy distribution:
$p(x)=\frac{1}{\pi(1+x^{2})}$ (22)
which has undefined mean and variance due to its extremely heavy tails.
Figure 38: In-context density estimation trajectories for t-distribution targets. Top row: 2D InPCA embeddings of DE trajectories for Mistral-7b-v0.3. Middle row: 2D InPCA embeddings of DE trajectories for LLaMA-2 13b. Bottom row: Corresponding ground truth distributions, with $v=0.1$, $1$, and $2$. These 2D embeddings capture 84% of pairwise Hellinger distances between probability distributions.
LLaMA and Mistral trace out similar in-context DE trajectories, as shown in Figure 38. Both models eventually converge to the ground truth distribution. However, unlike the cases with Gaussian and uniform targets, the trajectories for learning t-distributions are more erratic. This is likely due to the fact that t-distribution with small $v$ has large variance (diverging if $v\leq 1$), where the thick tails have non-trivial probability mass. This in turn makes extreme samples occur more frequently, causing instability in any kernel-like density estimators.
A.11 Additional notes on Bayesian Histogram
Figure 39: Density estimation trajectories for Bayesian histogram at different uniform bias $\alpha$. Top row: 2D InPCA embeddings of DE trajectories for Mistral-7b-v0.3. Bottom row: Corresponding ground truth distributions. These 2D embeddings capture 87% of pairwise Hellinger distances between probability distributions.
The Bayesian histogram with uniform prior starts near ignorance by construction, and gradually converges to ground truth distribution with increasing data. Its DE trajectories serve as informative visual comparisons against kernel-like methods. This work does not focus on Bayesian histogram per se. However, for completion, we briefly visualize the learning trajectories of Bayesian histograms at different levels of prior bias, $\alpha$, as defined in Equation 5. This results are shown in Figure 39
Unburdened by a strong prior, Bayesian histograms with smaller $\alpha$ converge faster to the ground truth. However, its learning trajectories are also more erratic. This is because in the absence of a strong regularization, any new data point would cause significant update in the estimated distribution (height of the corresponding bin). This effect is especially pronounced for high-variance distribution, where extreme data occur more often.
Bayesian histograms with larger $\alpha$, on the other hand, follow the geodesic more conservatively in smooth trajectories. However, they also take more data to converge to the true distribution. Note that unlike the other visualizations in this work, each trajectory in Figure 39 consists of $n=500$ data points, instead of $200$.
A.12 Uncertainty quantification for Bespoke KDE parameters
In this section, we provide information-theoretic justifications for our estimated uncertainty of the fitted KDE parameters, outlined in Fig. 5. On a high level, our method is based on the observation that the Hessian matrix at optimal fit is approximately equal to the Fisher information matrix (Kullback, 1997).
Let $P(x)$ denote an LLM’s estimated PDF from data, which we treat as the ground truth to fit to. Let $Q_{\theta}(x)$ denote the PDF estimated by parametrized KDE with parameter vector $\theta=(h,s)$. Technically, both $P(x)$ and $Q_{\theta}(x)$ are conditioned on the set of data observed in-context, $\{X_{i}\}_{i=1}^{n}$. For simplicity, we drop this conditioning in our notation.
During the fitting process, the loss function is defined as the KL-divergence between $P(x)$ and $Q_{\theta}(x)$:
$D_{\text{KL}}(P\|Q_{\theta})=\int P(x)\log\frac{P(x)}{Q_{\theta}(x)}\,\textup{d}x$ (23)
The second derivative (Hessian) of the loss function with respect to $\theta$ is:
[…]
The Fisher Information Matrix (FIM) is defined as:
[…]
Therefore, at the optimal fit $\theta^{*}$, if $Q_{\theta}(x)\approx P(x)$, then the Hessian of the KL divergence provides a good approximation to the FIM:
$H_{\text{KL}}(\theta^{*})=\nabla_{\theta}^{2}D_{\text{KL}}(P\|Q_{\theta})\approx I(\theta^{*})$
The covariance matrix of $\theta$ can then be estimated as the inverse of the FIM, a relationship known as the Cramér-Rao bound (Thomas & Joy, 2006). Under regularity conditions and asymptotic normality, this bound becomes an equality:
$\operatorname{Cov}(\theta)=I(\theta^{*})\approx\left(\nabla_{\theta}^{2}D_{\text{KL}}(P\|Q_{\theta})\right)^{-1}$
Said another way, the uncertainty in the fitted kernel parameters, such as bandwidth (h) and shape (s), can be estimated from the inverted Hessian matrix of KL-divergence loss function.
A.13 KDE with Silverman bandwidth schedule
@@
with $C=1$.However, as noted in Appendix A.7.1, in practice one often estimates the pre-coefficient from data, such as Silverman’s rule of thumb (Equation 17). In this section, we replicate Figures 3 and 4 from Section 4, and use Silverman’s rule-of-thumb bandwidth instead of $C=1$. As shown in Figures 40 and 41, even with Silverman’s pre-coefficient, Gaussian KDE still shows a clear bias towards the Gaussian submanifold.